# 4D Scaffold Gaussian Splatting for Efficient Dynamic Scene Reconstruction

### v1

@@ 3.4 Training and implementation details
K-Planes [24] 1 31.22 0.947 0.090 309 MB 0.13
MixVoxels-L [28] 1 30.81 0.933 0.095 512 MB 0.93
HyperReel [1] 1 2 31.10 0.927 0.099 1362 MB 1.04
4DGaussian [31] 30.71 0.935 0.056 59 MB 51.9
E-D3DGS [2] 31.42 0.945 0.037 137 MB 19.9
STG [14] 2 3 31.96 0.948 0.046 173 MB 125.9
4DGS [35] 32.14 0.947 0.047 6194 MB 61.4
Ours 32.03 0.947 0.041 134 MB 129.9
Training objective.   Both learnable anchor parameters and MLP weights are jointly optimized through the rendering loss. We employ $\mathcal{L}_{1}$ with SSIM loss $\mathcal{L}_{\text{SSIM}}$ and volume regularization $\mathcal{L}_{\text{vol}}$ following the 3D Scaffold-GS [16]. The full training objective with weighting coefficients $\lambda_{\text{SSIM}}=0.2$ and $\lambda_{\text{vol}}=0.01$ is
@@
Implementation details.   Similar to 4DGS, our model is trained for 120K iterations with a single batch, which takes approximately 4 hours on a single NVIDIA A6000 GPU. We set $\beta=2$ and $\gamma=1$ for our modified temporal opacity model and temporal coverage-aware densification. For 4D voxel grid size, We use 0.001 as spatial grid size and use frame interval time as temporal grid size for all scenes. After training, we prune the unnecessary anchor points of which all Gaussians have opacity near zero. To improve rendering speed, we cache the time-invariant and view-invariant outputs of MLPs. We provide more implementation details in the supplements.
4 Experiments
@@
Datasets.   We evaluate our method on two representative real-world datasets: Neural 3D Video dataset (N3DV) [12] and Technicolor dataset [23].
@@
Baselines.   We choose the following state-of-the-art competitors: 4DGaussian [31] and E-D3DGS [2] representing deformation-based methods, and 4DGS [35] and STG [14] representing 4D Gaussians. We reproduce them using their official code to report their performance. We also include representatives of NeRF-based methods: K-Planes [24], MixVoxels [28], and HyperReel [1]. Results of these methods are brought from [2].
Metrics.   To assess the reconstruction quality, we measure peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and LPIPS [38] of the rendered images. We employ AlexNet [9] to compute LPIPS following the previous papers. To evaluate storage efficiency, we calculate the total size of output files, including MLP weights and parameters of Gaussians or anchors. We measure frames-per-second (FPS) by recording the wall-clock rendering time on the same machine, with a single NVIDIA A6000 GPU.
Table 2: Quantitative results on Technicolor dataset. 1 refers to the values reported in [14]. 2 only includes the Painter for storage and FPS [2]. 3 uses COLMAP [25] for all 50 frames.
@@
DyNeRF [12] 1 31.80 - 0.140 30 MB 0.02
HyperReel [1] 2 32.32 0.899 0.118 289 MB 0.91
4DGaussian [31] 29.62 0.844 0.176 51 MB 34.8
E-D3DGS [2] 33.24 0.907 0.100 77 MB 36.5
STG [14] 2 3 33.40 0.918 0.082 55 MB 112.7
4DGS [35] 33.35 0.910 0.095 9106 MB 43.3
Ours 33.97 0.923 0.074 98 MB 124.5
Figure 5: Qualitative comparisons on the Neural 3D Video dataset. This figure provides qualitative comparisons on the Neural 3D Video dataset. White boxes indicate under-reconstructed areas. Our method achieves high fidelity in both dynamic and static regions.
4.1 Reconstruction quality
N3DV.   Table 1 presents the quantitative results on the N3DV dataset. Our method outperforms all NeRF baselines in both visual quality and rendering speed. It also achieves competitive results in visual quality compared to state-of-the-art Gaussian Splatting-based methods, and offers the fastest rendering speed across all methods. Remarkably, our method significantly reduces storage costs by 97.8×97.8$\times$97.8 × compared to 4DGS [35], while achieving comparable or better performance in visual quality and rendering speed. It is worth noting that STG requires COLMAP for all frames, and train 6 models to obtain its results. We provide further comparisons with STG in the supplements.
To further assess the performance, we deliver the qualitative comparisons with Gaussian Splatting baselines in Figure 5. Our method successfully reconstructs both static and dynamic regions. In contrast, other baselines often struggle to reconstruct dynamic regions (see the boxed regions). Deformation-based methods (4DGaussians, E-D3DGS) tend to produce blurry results, and the results of 4D Gaussian-based methods (4DGS, STG) often contain artifacts in dynamic regions. By effectively modeling dynamic regions with fewer anchors, our method delivers high-quality results in an efficient manner. Please refer to the supplements for more visual comparisons, including uncropped results and videos.
Figure 6: Qualitative comparisons on the Technicolor dataset. This figure provides qualitative comparisons on the Technicolor dataset. White boxes indicate under-reconstructed areas. Our method achieves high fidelity in both dynamic and static regions.
Technicolor.   We deliver the quantitative comparisons on the Technicolor dataset in Table 2. Our method surpasses all baselines in visual quality and rendering speed. We also achieve 85.1×85.1$\times$85.1 × storage reduction and $\times2$ improvement in FPS, compared to 4DGS. Although deformation-based methods offer compact storage usage, these methods require expensive computations for Gaussian deformations. This leads to the low FPS of these methods.
Figure 6 presents the qualitative comparisons with Gaussian-based methods on the Technicolor dataset. Our method produces sharp and high-fidelity rendering results in both static and dynamic regions. On the other hand, 4DGaussians, E-D3DGS and STG show blurry results in dynamic regions, such as the boxed part in the Theater scene. E-D3DGS and STG often fail to reconstruct several details, e.g. the eye in the Theater scene.
4.2 Analysis and ablation study
To provide deeper insights of our main features, we conduct the ablations and analysis on the N3DV dataset. Please refer to the supplements for more detailed analysis and ablations on each component.
Effects of each feature on quality and storage.   We verify the effectiveness of our proposals through the ablations. We begin with the baseline model (Baseline) that does not employ none of our main components. Then we gradually add each component to the baseline and report its results.
As shown in Table 3, our model with all components achieves the best results in visual quality with reasonable storage. Adding the temporal coverage-aware anchor growing operation improves visual quality but increases the anchors. The neural velocity reduces the number of anchors but brings degradation in visual quality. Finally, modifying the temporal opacity achieves the highest visual quality with fewer anchors. These results support that our anchor growing strategy successfully allocates new anchors to under-reconstructed dynamic regions, improving visual quality. Appropriate modeling of neural Gaussians reduces the number of anchors without sacrificing performance.
@@ 4.2 Analysis and ablation study
Table 3: Ablation results on the main components. The abbreviations are as follows: “TA” refers to the temporal coverage-aware anchor growing, “NV” to the neural velocity and “MT” to the modified temporal opacity.
@@
31.20 0.053 207K 83 MB
31.74 0.043  [^8] K 301 MB
31.11 0.042 576K 195 MB
31.82 0.043 793k 268 MB
32.03 0.041 440K 149 MB
Understanding anchor growing.   To provide an in-depth understanding of the proposed anchor growing strategy, we present visual comparisons with previous anchor growing operations. We visualize the accumulated gradients $\nabla_{g}$ of both methods during training. We also compare resulting anchors and rendered images after training.
Figure 7 presents visual comparisons with the previous anchor growing operation. The previous method fails to accurately collect gradients in dynamic regions due to their short appearing periods. This leads to a lack of anchors in under-reconstructed regions, resulting in degraded visual quality. In contrast, our method successfully accumulates higher gradients in these dynamic regions (see the red-boxed parts in Figure 7). Consequently, the anchors generated by our method more accurately represent dynamic regions at each timestep. These results demonstrate that our anchor-growing strategy effectively captures under-reconstructed dynamic regions, playing a critical role in achieving high-quality reconstruction.
Figure 8: Temporal opacity analysis. We visualize Gaussians that contributing to the rendering of a sample pixel in the sear_steak scene from the N3DV dataset. For clarity, we filter out Gaussians distant from the frying pan based on depth and image-plane probability. Our modified temporal opacity better aligns with real-world dynamics, allowing a single Gaussian to represent the element.
Understanding temporal opacity.   We explore the effects of the temporal opacity by observing actual distributions. Specifically, we choose an example pixel in sear_steak, which contains an object that appears in a certain period (the frying pan). We visualize Gaussians that primarily contribute to render this pixel. We compare our temporal opacity with the previous one based on a univariate Gaussian function.
As shown in Figure 8, the resulting Gaussian with our modified temporal opacity better fits the actual distributions of the object. This enables our method to represent a scene element with a single Gaussian. On the other hand, previous univariate Gaussian is not well-aligned to real-world dynamics, resulting in multiple Gaussians representing the same element. As a result, our temporal opacity reduces the number of anchors while retaining visual quality, which leads to compact storage usage.
5 Conclusion
We proposed the 4D anchor-based method for dynamic scene reconstruction. Our anchor growing and neural 4D Gaussians modeling capture dynamic regions with an efficient number of anchors. Experimental results support the validity of our method, achieving compelling visual quality and FPS while significantly reducing storage consumption.
Limitations and future work.   Since our method is currently designed for multi-view video datasets, applying it on monocular videos can introduce additional challenges. Same as other methods, our method still suffer from reconstructing elements that appear very shortly (1 or 2 frames). Resolving this challenge can be an interesting future work.
Supplementary Material
Figure S1: Comparison between 300-frame STG and Ours. The single STG model trained on 300 frames of the N3DV dataset shows lower quality in the dynamic regions compared to Ours. The white boxes highlight under-reconstruction areas.
@@ 5 Conclusion
[…]
STG 6 (50 frames) 6$\times$30K 31.96 0.046
1 (300 frames) 60K 31.25 0.053
Ours 1 (300 frames) 120K 32.03 0.041
Appendix A Additional comparison with STG
As mentioned in the main paper, it is difficult to fairly compare with Spacetime Gaussians (STG) [14] due to the different experimental setups. Unlike our method, STG utilizes per-frame COLMAP to obtain the initial points and divide long video sequences into 50-frame chunks, training separate models for each chunk. This results in six separate models to report its results on N3DV.
We additionally deliver the results of STG on N3DV, under the same experimental setup in Figure S1 and Table S1. While STG is effective on shorter video sequences, it often fails to accurately reconstruct dynamic regions on longer sequences. On the other hand, our method effectively represents dynamic regions on longer sequences. As a result, our method achieves better visual quality than STG under the same experimental setups.
Appendix B Modified temporal opacity with 4DGS.
We further validate the effectiveness of our modified temporal opacity by applying it to 4DGS [35]. Consistent with the main paper, we visualize temporal opacity values at the same points within sear_steak scene of the N3DV dataset. $\beta=8$ is used for the generalized Gaussian equation.
As shown in Figure S2, the previous univariate Gaussian function struggles to represent elements that suddenly appear and disappear. This limitation results in five Gaussians to model the same point. On the other hand, our modified temporal opacity effectively represents these dynamic changes, enabling the representation with only one Gaussian. Consequently, 4DGS with the modified temporal opacity achieves a storage reduction from 4.74 GB to 3.82 GB.
Figure S2: Impact of modified temporal opacity with 4DGS. We apply our modified temporal opacity to 4DGS and analyze its behavior on the sear_steak scene from N3DV. The total temporal opacity distribution for the frying pan (blue point) is low at $t=20$ and $t=90$ while peaking at $t=60$, with rapid transitions in between. Each plot represents the temporal opacity of an individual Gaussian. Unlike conventional methods relying on Gaussian mixtures, our generalized Gaussian effectively captures this behavior using a single Gaussian component.
Appendix C Additional ablation study
Modeling 4D Gaussians without anchors.   To analyze the effectiveness of our anchor-based reconstruction pipeline, we evaluate the performance of our model without the scaffold structure. Specifically, we replace all MLP outputs with explicit learnable Gaussian parameters. As shown in Table S2, removing the scaffold structure leads to a degradation in visual quality. The structured anchor representations offer more stable training and anchor growing, which significantly improves the performance of our method.
Polynomial Gaussian trajectory.   Our neural velocity represents the local dynamics of Gaussians as linear movements. Despite its simplicity, these linear movements sufficiently represent dynamics as 3D scene flows. We explore the effectiveness of the neural velocity by replacing it with a polynomial trajectory proposed in the previous method [14]. We use $n_{p}=3,n_{q}=1$ as they use for their full model. The velocity MLP outputs coefficients of the polynomial equation instead of the neural 3D velocity. As shown in Table S2, this approach results in decreased performance and increased storage. Since the polynomial trajectory requires 12 parameters per Gaussian, a total of 120 parameters ($K=10$) as MLP outputs, it often complicates the training of our lightweight MLP and leads to degradation.
Table S2: Additional model ablations. Quantitative results on different model designs on the N3DV dataset. Our full method utilizes scaffold structure and neural velocity.
[…]
velocity 30.11 0.060 393 MB
polynomial 31.68 0.046 323 MB
velocity 32.03 0.041 134 MB
Appendix D Effects of hyperparameters
@@
Temporal coverage-aware anchor growing.   The parameter $\gamma$ in Equation (5) controls the influence of temporal scale on densification. As $\gamma$ increases, densification becomes more concentrated in dynamic regions, improving reconstruction quality in these areas while reducing quality in static regions, as demonstrated in Figure S3. In scenes with distant background regions, such as flame_salmon, larger $\gamma$ values result in less densification in the static background, leading to a decline in reconstruction quality. Additionally, excessively high $\gamma$ values result in over-densification in dynamic regions, increasing the number of anchors. We empirically observe that $\gamma=1$ provides a balanced reconstruction between static and dynamic regions.
Modified temporal opacity.   The parameter $\beta$ determines the steepness of our generalized Gaussian curve. Higher values of $\beta$ result in steeper slopes, enabling shapes to be represented with fewer Gaussian components compared to a typical mixture of univariate Gaussians. As shown in Figure S4, increasing $\beta$ reduces the number of required anchor points while maintaining image quality. However, excessively high $\beta$ values (e.g., above 8) cause instability during training.
Appendix E Visualization of anchor growing
@@
Appendix F Additional method detail
@@
The Opacity MLP outputs the time-invariant opacity $\alpha$ of the Gaussians. The Shape MLP produces the quaternion q and scaling s for covariance calculation, along with the temporal scale $\sigma$. The Color MLP takes the direction d concatenated with the anchor feature as input and outputs the view-dependent color c. The Velocity MLP computes the 3D velocity of the Gaussians. $\tanh$ for opacity $\alpha$, $\exp$ for scaling s and temporal scale $\sigma$, and sigmoid for color c are applied as activation function.
Implmentation detail.   The temporal grid size is set to 0.0333 for the N3DV dataset and 0.02 for the Technicolor dataset. For initial points, we downsample it to fewer than 100,000 points. All other hyperparameters, including the learning rate and learning decay schedule, follow those of [16]. Components such as the feature bank, LOD, and appearance features in [16] are excluded.
Appendix G More results
We report the quantitative results in Table S3-S4 and qualitative results in Figure S6-S7 of each scene of the N3DV and the Technicolor datasets, respectively. We also provide video quality comparisons in HTML format. Please check the index.html file.
Table S3: Per-scene quantitative results on the N3DV dataset.
Model Metric coffee_martini cook_spinach cut_roasted_beef

### v2

@@ Introduction
4DGaussian (2023) 25.33 0.833 0.166 - 30.71 0.935 0.056 192 K 51.9 57 MB 50 m
E-D3DGS (2024) 26.92 0.884 0.112 - 32.04 0.951 0.034 180 K 74.5 66 MB 1 h 52 m
Grid4D (2024) 26.65 0.877 0.129 - 31.92 0.949 0.039 202 K 127.1 48 MB 1 h 20 m
STG (2023) 25.84 0.860 0.127 44 K 31.24 0.941 0.051 434 K 125.9 63 MB 1 h 8 m
4DGS (2024) 27.65 0.907 0.075 3306 K 32.14 0.947 0.047 3333 K 61.4 6194 MB 9 h 30 m
4DGS (1GB) 26.70 0.877 0.123 557 K 31.59 0.943 0.052 581 K 152.7 1080 MB 3 h 37 m
Ex4DGS (2024a) 26.33 0.874 0.121 52 K 32.01 0.947 0.048 268 K 91.9 115 MB 1 h 6 m
Scaff-naive 24.79 0.811 0.199 206 K 31.21 0.942 0.053 732 K 159.7 83 MB 2 h 57 m
Ours-light 27.50 0.902 0.076 181 K 31.54 0.944 0.045 314 K 148.2 90 MB 2 h 51 m
Ours 28.86 0.927 0.054 533 K 32.03 0.947 0.041 775 K 129.9 149 MB 3 h 6 m
Table 1: Quantitative results on the N3DV dataset. 4DGS refers to the result with 1GB storage, for fair comparisons.
Figure 4: Qualitative comparisons on the N3DV dataset. Our method achieves high-fidelity in both static and dynamic regions.
Training objective.  Both learnable anchor parameters and MLP weights are jointly optimized through the rendering loss. We employ $\mathcal{L}_{1}$ with SSIM loss $\mathcal{L}_{\text{SSIM}}$ and volume regularization $\mathcal{L}_{\text{vol}}$ following the 3D Scaffold-GS (Lu et al. 2024a). The full training objective with weighting coefficients $\lambda_{\text{SSIM}}=0.2$ and $\lambda_{\text{vol}}=0.01$ is
@@
Implementation details.  Similar to 4DGS, our model is trained for 120K iterations with a single batch, which takes approximately 3 hours on a single NVIDIA A6000 GPU. We set $\beta=2$ and $\gamma=1$ for our modified temporal opacity model and dynamic-aware densification. For 4D voxel grid size, we use 0.001 as the spatial grid size and use the frame interval time as the temporal grid size for all scenes. We set $K=10$ following (Lu et al. 2024a). After training, we prune the invalid anchor points of which all Gaussians have negative opacity. To improve inference speed, we cache the time-invariant and view-invariant outputs of MLPs. We provide more implementation details in the supplements.
Dynamic region Full region Computational cost
[…]
4DGaussian (2023) 23.99 0.709 0.280 - 29.62 0.843 0.176 268 K 23.4 72 MB 31 m
E-D3DGS (2024) 29.39 0.886 0.148 - 33.38 0.907 0.100 212 K 60.8 77 MB 2 h 34 m
STG (2023) 28.35 0.869 0.174 71 K 33.33 0.912 0.096 132 K 149.0 30 MB 1 h 24 m
4DGS (2024) 31.91 0.930 0.093 5597 K 33.30 0.910 0.095 5759 K 141.5 10699 MB 6 h 11 m
4DGS (1GB) 29.18 0.890 0.144 538 K 28.72 0.856 0.166 591 K 156.6 1098 MB 3 h 36 m
C3DGS (2024c) 27.21 0.843 0.209 87 K 32.52 0.901 0.110 122 K 178.7 18 MB 1 h 9 m
Ex4DGS (2024a) 30.58 0.918 0.109 120 K 33.40 0.914 0.088 426 K 78.4 140 MB 1 h 53 m
Scaff-naive 28.00 0.853 0.195 721 K 33.34 0.915 0.081 2056 K 127.0 129 MB 1 h 59 m
Ours-light 30.95 0.916 0.117 186 K 33.97 0.923  0.074 480 K 131.4 108 MB 1 h 40 m
Ours 31.94 0.932 0.093 1165 K 34.11 0.925 0.072 1272 K 110.8 278 MB 2 h 30 m
Table 2: Quantitative results on Technicolor dataset. For fair comparison, 4DGS result of using 1GB storage is reported.
Experiments
@@
Datasets.  We evaluate our method on two representative real-world datasets, which are Neural 3D Video (N3DV) (Li et al. 2022) and Technicolor (Sabater et al. 2017).
@@
Baselines.  We choose the following state-of-the-art competitors: 4DGaussian (Wu et al. 2023), E-D3DGS (Bae et al. 2024) and Grid4D (Jiawei et al. 2024) representing deformation-based methods, and 4DGS (Yang et al. 2024), STG (Li et al. 2023), C3DGS (Lee et al. 2024c) and Ex4DGS (Lee et al. 2024a) representing 4D Gaussians. We reproduce them using the official code to report their performance. We also introduce Scaff-naive as an additional baseline. This is an anchor-based model without our neural Gaussian design and the anchor growing, which instead adopts linear motion and the temporal opacity modeling used in (Yang et al. 2024). We present a variant of our model, Ours-light, which employs a larger voxel size to achieve a more reduced storage footprint. For fair comparisons, we report STG trained using 300 frames and initial SfM points derived from the first frames, same as other methods.
Metrics.  To assess the reconstruction quality, we measure peak signal-to-noise ratio (PSNR), structural similarity index (SSIM), and LPIPS (Zhang et al. 2018) of the rendered images. To compute LPIPS, we follow (Bae et al. 2024) and employ AlexNet (Krizhevsky, Sutskever, and Hinton 2012). To measure the visual quality in dynamic regions, we use a combined mask of Global-Median and Temporal-Difference (Li et al. 2022), binarized with a threshold of 50. We assess storage efficiency by calculating the total size of the output files, including MLP weights, as well as Gaussian and anchor parameters. In addition to storage size, we also report the number of Gaussians in both dynamic and full regions. FPS and training time are measured on a single NVIDIA A6000 GPU, using the flame_salmon and Fabien scenes from N3DV and Technicolor, respectively.
Experimental Results
N3DV.  Table 1 presents the quantitative results on the N3DV dataset. Our model outperforms all baselines in terms of visual quality in dynamic regions, while also delivering competitive results in full regions. While our method employs more Gaussians to achieve high-quality results, it maintains efficient storage overhead and FPS, thanks to the anchor-based compression scheme. Our storage-efficient variant (Ours-light) achieves the second-best storage efficiency among 4D Gaussian methods, while also surpassing other efficient baselines in visual quality. Scaff-naive yields degraded results, indicating that the naïve extension of the scaffolding is insufficient to accurately reconstruct dynamic 3D scenes. 4DGS often models static components through multiple dynamic Gaussians, resulting in an excessive number of Gaussians and storage costs. Although our method does not leverage additional techniques to model far-background areas (Li et al. 2023; Yang et al. 2024), it still exhibits plausible reconstruction quality in both static and full regions.
The effectiveness of our method is also demonstrated in qualitative results. As shown in Figure 4, our model effectively represents dynamic regions with a sufficient number of Gaussians, thereby achieving high-quality dynamic reconstruction. In contrast, other efficient baselines struggle to capture complex scene dynamics (see the boxed regions). Please refer to the supplements for more visual comparisons, including uncropped results and videos.
Figure 5: Quality-storage trade-off. The results on the flame_steak scene of N3DV, along with the number of Gaussians, are reported. We increase the number of employed Gaussians of our model by adjusting the anchor grid size.
@@
Polynomial Ours 27.82 0.070 257 MB
Linear 4DGS 27.93 0.065 195 MB
Linear Ex4DGS 28.34 0.056 413 MB
Linear Ours 25.77 0.153 180 MB
Linear Ours 29.57 0.050 149 MB
@@
Table 3: Ablation results on the main components. “DA” refers to the dynamic-aware anchor growing, “Motion” and “Opacity” for each modeling of the neural Gaussian design.
Technicolor.  We deliver the quantitative comparisons on the Technicolor dataset in Table 2. Our model achieves state-of-the-art performance across all visual quality evaluation metrics. This is further exemplified by our storage-efficient model, Ours-light. Despite its compact size of approximately 100 MB, Ours-light delivers competitive visual quality on par with 4DGS, highlighting the effectiveness of the proposed anchor-based scheme. Qualitative comparisons on the Technicolor dataset are provided in the supplements.
Analysis and Ablation Study
To provide deeper insights of our main features, we conduct the ablations and analysis in dynamic regions on N3DV.
Quality-storage trade-off.  We first analyze the quality-storage trade-off of our model by gradually increasing the number of employed Gaussians. This is controlled by adjusting the voxel size of the anchor grid. As presented in Figure 5, the performance of our model is improved with more Gaussians, confirming the necessity of using a sufficient quantity. Notably, our method outperforms other storage-efficient baselines when using similar or smaller storage costs. Thanks to the anchor-based scheme, our method employs $3.37\times$ or $7.0\times$ more Gaussians and achieves higher visual quality, yet still maintains a lower storage overhead.
Understanding the anchor growing.  We then ablate the proposed anchor growing strategy. As reported at the bottom of Table 3, the dynamic-aware anchor growing leads to noticeable improvement in visual quality. By properly allocating new anchors to under-reconstructed dynamic regions, our model accurately represents dynamic scenes with practical storage overhead.
To further explore the effects of the anchor growing strategy, we provide visual comparisons in Figure 6. The previous method fails to accurately collect gradients in dynamic regions due to their short appearing periods. This leads to a lack of anchors in dynamic regions, resulting in degraded visual quality. In contrast, our method successfully accumulates higher gradients in these dynamic regions (see the red-boxed parts in Figure 6). Consequently, the anchors generated by our method more accurately represent dynamic regions at each timestep. These results demonstrate that our anchor-growing strategy effectively captures under-reconstructed dynamic regions, playing a critical role in achieving high-quality reconstruction.
Effects of neural Gaussian design.  We first evaluate our neural design by replacing each modeling with the formulation in previous works. Specifically, we replace the linear motion with the polynomial trajectory (Li et al. 2023), and replace our modified temporal opacity with the one used in 4DGS (Yang et al. 2024) and Ex4DGS (Lee et al. 2024a). As shown in Table 3, our model with the proposed components achieves the best results. Our compact parametrization enhances the expressiveness of each Gaussian with minimal parameters, improving both visual quality and efficiency.
Conclusion
Our framework employs a sufficient number of Gaussians to capture complex dynamic regions, while addressing the storage overhead through the anchor-based compression. The dynamic-aware anchor growing and the neural Gaussian design lead to a substantial improvement. Experimental results support the validity of our method, achieving state-of-the-art visual quality and practical storage costs.
Limitations and future work.  Since our method is currently designed for multi-view video datasets, applying it to monocular videos can introduce additional challenges. Like other methods, our method still suffers from reconstructing elements that appear very shortly (1 or 2 frames). Resolving this challenge can be an interesting future work.
Appendix
Additional Comparison with STG
Unlike our approach, which operates from sparse initial points, STG’s (Li et al. 2023) original setting utilizes dense SfM point clouds from all timesteps and train the scene in 50-frame sequences. This original setting requires significant preprocessing for COLMAP (Schönberger and Frahm 2016) calculation and less densification process during training. Table S1 and Figure S1 show a qualitative and quantitative comparison between STG and our method, using the same number of iterations. While STG achieves higher scores when operating in its original dense setting, our method still demonstrates a superior score in visual quality. This indicates that our method successfully reconstructs fine-grained details that were omitted by the initial SfM points.
Figure S1: Comparison between 50-frame STG and Ours. The single STG model trained on 300 frames of the N3DV dataset shows lower quality in the dynamic regions compared to Ours. The white boxes highlight under-reconstruction areas.
@@ Conclusion
[…]
Ours 1 (300 frames) 120K 32.03 0.041
@@
Understanding Temporal Opacity
We explore the effects of the temporal opacity by observing actual distributions. Specifically, we choose an example pixel in sear_steak, which contains an object that appears in a certain period (the frying pan). We visualize Gaussians that primarily contribute to render this pixel. We compare our temporal opacity with the previous one based on a univariate Gaussian function.
As shown in Figure S2, the resulting Gaussian with our modified temporal opacity better fits the actual distributions of the object. This enables our method to represent a scene element with a single Gaussian. On the other hand, previous univariate Gaussian is not well-aligned to real-world dynamics, resulting in multiple Gaussians representing the same element. As a result, our temporal opacity reduces the number of anchors while retaining visual quality, which leads to compact storage usage.
Figure S2: Temporal opacity analysis. We visualize Gaussians contributing to the rendering of a sample pixel in the sear_steak scene from the N3DV dataset. For clarity, we filter out Gaussians distant from the frying pan based on depth and image-plane probability Our modified temporal opacity better aligns with real-world dynamics, enabling a single Gaussian to represent the element.
Effects of Hyperparameters
@@
$K=3$ $K=5$ $K=8$ $K=10$
PSNR 27.40 28.63 29.09 29.20
#Anchors 579K 585K 494K 476K
Table S2: Results with various $K$.
@@
Number of Gaussians per anchor.  The parameter $K$ determines the number of Gaussians generated from an anchor feature. We observed changes in image quality and the number of anchors as $K$ varied. As $K$ increased, image quality improved while the number of anchors decreased. This is because a larger $K$ allows the same spatial region to be represented with higher density.
Dynamic-aware anchor growing.  The parameter $\gamma$ in Equation (9) controls the influence of temporal scale on densification. As $\gamma$ increases, densification becomes more concentrated in dynamic regions, improving reconstruction quality in these areas while reducing quality in static regions, as demonstrated in Figure S3. In scenes with distant background regions, such as flame_salmon, larger $\gamma$ values result in less densification in the static background, leading to a decline in reconstruction quality. Additionally, excessively high $\gamma$ values result in over-densification in dynamic regions, increasing the number of anchors. We empirically observe that $\gamma=1$ provides a balanced reconstruction between static and dynamic regions.
Temporal opacity.  The parameter $\beta$ determines the steepness of our generalized Gaussian curve. Higher values of $\beta$ result in steeper slopes, enabling shapes to be represented with fewer Gaussian components compared to a typical mixture of univariate Gaussians. As shown in Figure S4, increasing $\beta$ reduces the number of required anchor points while maintaining image quality. However, excessively high $\beta$ values (e.g., above 8) cause instability during training.
Visualization of Anchor Growing
@@
Additional Method Details
@@
The Opacity MLP outputs the time-invariant opacity $\rho$ of the Gaussians. The Shape MLP produces the quaternion q and scaling s for covariance calculation, along with the temporal scale $\sigma$. The Color MLP takes the direction d concatenated with the anchor feature as input and outputs the view-dependent color c. The Velocity MLP computes the 3D velocity u of the Gaussians. $\tanh$ for opacity $\rho$, $\exp$ for scaling s and temporal scale $\sigma$, and sigmoid for color c are applied as activation function.
Additional implementation detail.  The temporal grid size is set to 0.0333 for the N3DV dataset and 0.02 for the Technicolor dataset. For initial points, we downsample it to fewer than 100,000 points. All other hyperparameters, including the learning rate and learning decay schedule, follow those of Scaffold-GS (Lu et al. 2024a). Components such as the feature bank, level of detail (LOD), and appearance features are excluded.
Results in Challenging Scenarios
PSNR SSIM Train_time Storage
4DGS 24.28 0.833 4 h 30 m 2359 MB
Ours 28.82 0.913 2 h 25 m 114 MB
Table S3: Results with limited views.
Few shots.  To evaluate the robustness of our method, we report the results in a 4 shot scenario, in which sparse 4 views are used for reconstruction. Results in a 4 shot scenario on the N3DV dataset are reported in Table S3. While our method was not specifically designed for few-shot scenarios, it exhibits a smaller performance drop compared to 4DGS.
More Results
We report the quantitative results in Table S4-S5 and qualitative results in Figure S7-S8 of each scene of the N3DV and the Technicolor datasets, respectively. We also provide video quality comparisons in HTML format. Please check the index.html file.
Figure S6: Qualitative comparisons on the Technicolor dataset.
Dynamic region Full region